### v1

# DataChef: Cooking Up Optimal Data Recipes for LLM Adaptation via Reinforcement Learning

Abstract
In the current landscape of Large Language Models (LLMs), the curation of large-scale, high-quality training data is a primary driver of model performance. A key lever is the data recipe, which comprises a data processing pipeline to transform raw sources into training corpora. Despite the growing use of LLMs to automate individual data processing steps, such as data synthesis and filtering, the overall design of data recipes remains largely manual and labor-intensive, requiring substantial human expertise and iteration. To bridge this gap, we formulate end-to-end data recipe generation for LLM adaptation. Given a target benchmark and a pool of available data sources, a model is required to output a complete data recipe that adapts a base LLM to the target task. We present DataChef-32B, which performs online reinforcement learning using a proxy reward that predicts downstream performance for candidate recipes. Across six held-out tasks, DataChef-32B produces practical recipes that reach comparable downstream performance to those curated by human experts. Notably, the recipe from DataChef-32B adapts Qwen3-1.7B-Base to the math domain, achieving 66.7 on AIME’25 and surpassing Qwen3-1.7B. This work sheds new light on automating LLM training and developing self-evolving AI systems.
DataChef: Cooking Up Optimal Data Recipes for LLM Adaptation
via Reinforcement Learning
Yicheng Chen1,2, Zerun Ma2, Xinchen Xie2, Yining Li2†, Kai Chen2† 1Fudan University 2Shanghai AI Laboratory Github: https://github.com/yichengchen24/DataChef
Figure 1: (a) Formulation. Given a task instruction, evaluation protocol, and raw data sources, a model is required to generate a data recipe, including an executable pipeline and the resulting training dataset, for LLM adaptation. (b) Main results. DataChef matches the performance of recipes from Gemini-3-Pro across six held-out tasks. See details in Sec. 4.2.
{NoHyper}footnotetext: Corresponding Author.
1 Introduction
The rapid advancement of Large Language Models (LLMs) DeepSeek-AI et al. (2025); OpenAI (2025) has underscored the central role of data in determining model capabilities, making a shift toward data-centric AI Jakubik et al. (2024). The composition and quality of training data emerge as decisive factors in shaping downstream performance. Practically, constructing effective training data requires a well-designed multi-stage pipeline that processes heterogeneous raw data through a sequence of operations, such as transformation, filtering, mixing, synthesis, and refinement, tailored to specific training goals or stages Yang et al. (2025); Cai et al. (2025). In this work, we formalize the concept of a data recipe, defined as the combination of the processing pipeline and the resulting training data.
In practice, constructing a data recipe typically involves substantial human expertise and manual effort, with human experts orchestrating a subset of data processing operations in a specific order Penedo et al. (2025); Gururajan et al. (2024). While LLMs are widely used in individual processing operations, such as data filtering Liu et al. (2024), data selection Zhang et al. (2025d), and data synthesis Mitra et al. (2024); Huang et al. (2024), they still follow a human-designed prompt or pattern to prepare data. Recent studies have explored automating data pipeline orchestration to reduce the reliance on manual effort. In particular, Data-Juicer Sandbox Chen et al. (2025a) proposes a Probe-Analyze-Refine workflow to identify the most impactful operators from a predefined operation pool, combine effective operations, and optimize data utilization through systematic experiments in data processing, model training, and evaluation with model performance as feedback. However, the continuous scaling of data and model sizes, coupled with the increasing complexity of processing operations, renders an exhaustive exploration of the combinatorial space of data recipes infeasible. Therefore, an essential question arises: can AI systems automatically generate a data recipe for training LLMs, including the orchestration of data pipelines and the implementation of each operation, in a cost-efficient way?
To bridge this gap, we introduce a new task: end-to-end data recipe generation for LLM adaptation. As shown in Fig 1(a), given a target benchmark and a pool of available data sources, the objective is to generate a complete data recipe by specifying the precise data processing pipeline to yield training data for adapting an LLM to the target task. Producing an effective data recipe requires strong reasoning abilities, as it involves analyzing heterogeneous data sources, applying domain-specific processing operations, and generating corresponding executable code. While recent studies have demonstrated the effectiveness of reinforcement learning in enhancing LLM reasoning abilities in complex domains, such as coding Zeng et al. (2025) and mathematical reasoning Yeo et al. (2025), applying this paradigm to our task poses two key challenges: (1) Data absence: end-to-end data recipe generation is a previously unexplored task, for which no curated datasets or standardized evaluation benchmarks exist. (2) Expensive and delayed supervision: while downstream performance naturally serves as the reward signal, it is impractical to directly incorporate downstream LLM training into an online reinforcement learning loop.
To address these challenges, we curate a comprehensive task pool comprising 31 widely used benchmarks across 10 distinct domains. These domains encompass reasoning-heavy fields, such as mathematics and coding, as well as knowledge-centric fields, such as finance, medicine, and the natural sciences. The pool is partitioned into 25 training tasks and 6 held-out evaluation tasks, with each task supported by 8–15 source training datasets. We further propose a data verifier that assesses training data quality directly without performing model training, providing a low-cost, instant reward signal in online RL. We systematically validate that the data verifier prediction correlates well with downstream model performance across domains. Leveraging cold-start fine-tuning and online reinforcement learning, we present DataChef-32B, an LLM specialized for generating optimal data recipes.
Through extensive evaluations, DataChef-32B demonstrates matching capability to Gemini-3-Pro on data recipe generation. Fig. 1(b) shows that DataChef-32B produces effective data recipes on 6 hold-out tasks, yielding high performance on downstream benchmarks. The generated recipes outperform the best individual data source on most tasks. Notably, on the math and atmosphere domain, recipes from DataChef-32B adapt Qwen3-1.7B-Base to achieve 66.7 on AIME’25 AIME (2025) and 46.3 on ClimaQA Manivannan et al. (2025), respectively, surpassing Qwen3-1.7B with industry-level post-training on expert-curated data recipes.
In summary, our contributions are as follows:
  1. $\bullet$
    We formulate a new task, end-to-end data recipe generation for LLM adaptation, requiring models to automatically generate data recipes from a benchmark and available data sources.
  2. $\bullet$
    We construct a large-scale and diverse data pool covering 19 domains, 31 benchmarks, and 257 datasets to support this task.
  3. $\bullet$
    We propose an efficient learning framework with a proxy reward that enables scalable online RL. Extensive experiments show that our DataChef-32B achieves performance comparable to that of top-tier proprietary models on the data recipe generation task.
2 Related Work
Data Pipelines. Many existing approaches rely on human experts to design individual data processing heuristics, including data mixing Liu et al. (2025b), data sampling Xu et al. (2024); Chen et al. (2025d), and data synthesis Chen et al. (2025c). General-purpose data processing frameworks Chen et al. (2024a); Park et al. (2025) provide standardized modules and scalable pipeline construction for large-scale data processing, and are adopted to curate large-scale, high-quality training data, such as FinWeb2 Penedo et al. (2025) for multilingual pre-training and Aloe Gururajan et al. (2024) for medical-domain fine-tuning. However, their efficiency remains constrained by the manual pipeline design and iterative trial-and-error on downstream tasks. Data-Juicer Sandbox Chen et al. (2025a) marking a step further towards automated data pipeline construction by employing a Probe-Analyze-Refine workflow to assess operator effectiveness, but still relies on feedback derived from downstream model training, which is time and computation-consuming. In contrast, our work aims to end-to-end generate data recipes from scratch.
LLM Agents for Data Science. LLM-based agent systems have emerged as powerful tools for automating data science workflows, including data analysis, modeling, and visualization. Most existing approaches Hollmann et al. (2023); Li et al. (2024b); Hong et al. (2025) rely on prompt-based approaches, where complex tasks are decomposed and solved according to heuristically designed workflows. AIDE Jiang et al. (2025) and SELA Chi et al. (2024) further adopt iterative exploration and refinement through trial-and-error execution. Yet such prompt-driven strategies remain largely static and are constrained by the inherent knowledge limitations of LLMs. To alleviate these limitations, some studies incorporate external knowledge via search-based methods, leveraging offline repositories such as Kaggle solutions and research papers Guo et al. (2024); Ou et al. (2025); Kulibaba et al. (2025) or online web search Nam et al. (2025). Another line of work Liu et al. (2025c); Zhang et al. (2025c) explores learning-based agents, where agents improve performance through interaction and experience. However, these methods are typically evaluated on well-defined Kaggle competitions Chan et al. (2025); Zhang et al. (2025b); Jing et al. (2025) with static datasets, and even with curated initial code. In this work, we address an open-ended setting, taking arbitrary tasks and available datasets as input and directly generating data recipes for LLM training.
Data Evaluation. Training and evaluating LLMs require significantly more computational resources, motivating the use of lightweight proxies to assess model performance Chen et al. (2025a). Existing data evaluation approaches Qin et al. (2024); Zhang et al. (2025a) can be broadly categorized into three groups. (1) Indicator-based methods Li et al. (2024a); Friedman and Dieng (2023) define handcrafted metrics to quantify properties such as diversity, complexity, and relevance. (2) Model-based methods Ge et al. (2024); Liu et al. (2024) train predictive models to estimate data quality. (3) LLM-as-a-Judge approaches Chen et al. (2024b) prompts powerful LLMs to evaluate data according to specific protocols. However, the correlation between data assessment scores and downstream model performance remains underexplored. Prior work typically validates evaluators by comparing specific data selections against baselines, rather than through systematic correlation analysis. To bridge this gap, we conduct a comprehensive study of representative assessment methods, evaluating their alignment with model performance across diverse fine-tuning tasks.
3 Methodology
In this section, we first formalize some core concepts and define the data recipe generation task in Sec 3.1. Then, we introduce the specific data pool constructed for this study in Sec 3.2. Finally, we present our learning framework in Sec 3.3.
Figure 2: Illustration of DataChef training framework. Given a task, a policy LLM generates a data recipe, which is executed to produce a training dataset. The Data Verifier then evaluates a sampled subset to provide a scalar reward, guiding the policy update via GRPO to optimize for data quality and executability.
3.1 Problem Formulation
The goal of our method is to automatically generate a data recipe given a specific task. We formulate a task as a triplet $T=(I,\tau,\mathcal{D})$, where $I$ is a natural language instruction, including description of the task requirement, along with meta-information of data sources and evaluation protocol, $\mathcal{D}$ denotes the set of available raw data sources, and $\tau$ is an evaluation metric that maps any model $\mathcal{M}$ to a scalar performance score $\tau(\mathcal{M})\in\mathbb{R}$. A data recipe is formulated as $r=(g,d)$, where $g\in\mathcal{G}$ is a data pipeline and $d=g(\mathcal{D})$ is the resulting training dataset. In our experiments, the data pipeline is implemented as Python scripts.
Let $\mathcal{M}_{\theta}$ denote a language model. We use $\theta_{d}$ to present the parameters fine-tuned on a dataset $d$. We aim to learn a policy $\pi_{\phi}(r\mid T)$ that generates data recipes to maximize the expected downstream performance of the trained model. Formally, the objective function is defined as:
$\mathcal{J}(\phi)=\mathbb{E}_{r\sim\pi_{\phi}(\cdot\mid T)}[\tau(\mathrm{LM}_{\theta_{d}})]$ (1)
3.2 Task Pool Construction
Seed Task Curation. As a previously unexplored task, data recipe generation lacks a canonical corpus. To bridge this gap, we construct a diverse task pool encompassing 19 heterogeneous domains, including reasoning, coding, and knowledge-intensive fields such as healthcare, finance, and natural science. For each domain, we select representative benchmarks (e.g., GSM8K and AIME’25 for mathematics), totaling 31 benchmarks. For each benchmark, we retrieve relevant candidate datasets from Hugging Face, prioritizing those with high community engagement (downloads and likes), yielding a repository of 257 distinct data sources. From this collection, we construct 25 seed tasks for training and reserve 6 held-out tasks (3 in-domain, 3 out-of-domain) for evaluation. Comprehensive details of the selected benchmarks are provided in Appx. A.1.
Training Task Augmentation. To facilitate robust policy learning, we expand the 25 seed tasks into a large-scale training set $\mathcal{T}_{\mathrm{train}}$. We employ a probabilistic sampling strategy where a benchmark $\tau$ is selected proportional to its source count $|\mathcal{D}|$, followed by uniform sampling of a subset $\mathcal{D}^{\prime}\subseteq\mathcal{D}$ to form a new instance $T^{\prime}=(I^{\prime},\tau,\mathcal{D}^{\prime})$. After deduplication, the expansion strategy yields 5K unique task instances.
3.3 End-to-end Data Recipe Generation
Framework Overview. As illustrated in Fig. 2, our framework optimizes the policy $\pi_{\phi}$ to generate high-quality data recipes. Given a task $T$, the policy generates a data pipeline $g$, which consists of a natural language plan for orchestrating data pipelines and its corresponding implementation as an executable code block. During training, the pipeline transforms raw data sources $\mathcal{D}$ into a training dataset $d$, which is then evaluated by the Data Verifier to guide policy updates via reinforcement learning. During inference, the data recipe is directly used for downstream model adaptation.
Cold-start Initialization. Training the policy from scratch using RL is non-trivial due to the low executability of data recipes, leading to sparse, high-variance rewards and ineffective exploration Shao et al. (2024); Liu et al. (2025c). To mitigate this, we employ a cold-start Supervised Fine-Tuning (SFT) phase. We observe that decoupling reasoning and coding yields superior inference-time performance (as discussed in Sec. 4.4). Therefore, we construct a high-quality demonstration set using a decoupled generation process: a strong reasoning model proposes plans, and a specialized coding model implements them. We filter these pairs by execution success and data quality, retaining only valid recipes. Initializing $\pi_{\phi}$ on this curated dataset equips the policy with a foundational capability for code generation, significantly stabilizing the subsequent RL phase.
Table 1: Main Results on six held-out tasks. We report the mean Data Verifier Score $\mathrm{DVS}_{\mathrm{avg}@32}$ and the Downstream Benchmark Score $\mathrm{DBS}$, where the Average column presents $\mathrm{DBS}$ as a normalized percentage relative to Sourcebest (100.0). Qwen3-Next $\oplus$ Kimi-K2 denotes a combination using Qwen3-Next-80B for reasoning and Kimi-K2-Instruct for coding. DataChef-32B achieves performance comparable to the closed-source Gemini-3-Pro and significantly outperforms other open-source baselines across all settings.
Indomain Tasks
Method PHYSICS AIME LiveCode Average
$\mathrm{DVS}_{\mathrm{avg}@32}\uparrow$ $\mathrm{DBS}\uparrow$ $\mathrm{DVS}_{\mathrm{avg}@32}\uparrow$ $\mathrm{DBS}\uparrow$ $\mathrm{DVS}_{\mathrm{avg}@32}\uparrow$ $\mathrm{DBS}\uparrow$ $\mathrm{DVS}_{\mathrm{avg}@32}\uparrow$ $\mathrm{DBS}\uparrow$
Sourceavg - 6.1 - 23.4 - 6.3 - 63.9
Sourcebest - 8.5 - 39.6 - 10.3 - 100.0
Expert - 20.8 - 33.3 - 25.7 - 123.3
Qwen3-32B 11.0 5.9 31.5 13.3 24.3 8.0 22.3 56.7
Kimi-K2 19.7 9.0 35.4 20.0 19.3 9.7 24.8 83.7
Qwen3-Next $\oplus$ Kimi-K2 48.7 8.9 78.3 23.3 39.2 7.4 55.4 78.6
Gemini-3-Pro 69.7 9.2 80.7 30.0 53.6 9.1 68.0 91.2
DataChef-32B 61.4 8.7 84.7 30.0 45.8 9.1 64.0 89.3
DataChef-32B (Oracle) - 10.4 - 66.7 - 10.3 - 130.3
Out-of-domain Tasks
Method ClimaQA OpenFin CHID Average
$\mathrm{DVS}_{\mathrm{avg}@32}\uparrow$ $\mathrm{DBS}\uparrow$ $\mathrm{DVS}_{\mathrm{avg}@32}\uparrow$ $\mathrm{DBS}\uparrow$ $\mathrm{DVS}_{\mathrm{avg}@32}\uparrow$ $\mathrm{DBS}\uparrow$ $\mathrm{DVS}_{\mathrm{avg}@32}\uparrow$ $\mathrm{DBS}\uparrow$
Sourceavg - 26.0 - 41.7 - 49.3 - 65.1
Sourcebest - 43.6 - 63.7 - 70.3 - 100.0
Expert - 44.2 - 73.4 - 59.8 - 101.0
Qwen3-32B 20.6 35.6 34.9 23.8 1.1 12.3 18.9 45.5
Kimi-K2 18.3 41.8 51.5 46.5 23.1 3.9 31.0 58.2
Qwen3-Next $\oplus$ Kimi-K2 41.5 42.6 54.7 64.0 8.6 4.1 34.9 68.0
Gemini-3-Pro 58.4 44.3 54.9 61.8 29.7 21.9 47.6 76.6
DataChef-32B 57.3 42.1 67.0 63.9 7.9 20.5 44.1 75.4
DataChef-32B (Oracle) - 46.3 - 67.1 - 45.7 - 92.2
Reward Modeling. Ideally, the reward signal would be the downstream performance $\tau(\mathcal{M}_{\theta_{d}})$. However, using this as an online reward is computationally prohibitive due to the cost of repeated model training and evaluation. Instead, we design a computationally efficient surrogate reward based on the quality of the generated dataset $d$. Inspired by rubrics-based rewards Gunjal et al. (2025), we employ a strong LLM as a Data Verifier to classify each instance $x\in d$ into one of five categories with assigned scalar scores $s(x)$:
  1. $\bullet$
    Invalid (0): Samples with missing essential information or severe repetition.
  2. $\bullet$
    Format Error (0): Samples violating explicit output format constraints.
  3. $\bullet$
    Incorrect (0): Samples containing factual errors or wrong answers.
  4. $\bullet$
    Task Mismatch ($0.4$): Valid samples that are semantically irrelevant to the target task $I$.
  5. $\bullet$
    Pass ($1.0$): High-quality samples that satisfy all criteria.
[…]
$\mathcal{R}(r)=\begin{cases}-\lambda_{\emptyset},&\text{if}\ d=\emptyset\ \text{(execution failure)},\\ -\lambda_{\mathrm{fmt}},&\text{if}\ d\ \text{{violates training format}},\\ \bar{s}(\hat{d}),&\text{{otherwise}},\end{cases}$ (2)
where $\lambda_{\emptyset}$ and $\lambda_{\mathrm{fmt}}$ are positive penalty coefficients. Please refer to Appx. A.2 for a detailed description of the category definitions used in the prompt.
Reinforcement Learning. We employ Group Relative Policy Optimization (GRPO) for policy optimization. For each task $T\sim\mathcal{T}_{\mathrm{train}}$, we sample a group of $G$ candidate data recipes $\{r_{i}\}_{i=1}^{G}$ from the current policy $\pi_{\phi_{\mathrm{old}}}$. The policy parameters are optimized by maximizing the following objective:
$\mathcal{J}(\phi)=\mathbb{E}\Bigg[\frac{1}{G}\sum_{i=1}^{G}\min\Big($ $\rho_{i}A_{i},\;\mathrm{clip}(\rho_{i},1-\epsilon,1+\epsilon)\,A_{i}\Big)$ (3)
$-\beta\,D_{\mathrm{KL}}\!\Big(\pi_{\phi}\,\|\,\pi_{\mathrm{ref}}\Big)\Bigg]$
where $\rho_{i}=\frac{\pi_{\phi}(r_{i}\mid T)}{\pi_{\phi_{\mathrm{old}}}(r_{i}\mid T)}$ is the importance ratio, $A_{i}=\frac{\mathcal{R}(r_{i})-\mu}{\sigma+\delta}$ is the group-relative advantage, $\epsilon$ is the clipping parameter, $\pi_{\mathrm{ref}}$ is a fixed reference policy, and $\beta$ controls KL regularization.
4 Experiments
4.1 Setups
Figure 3: Correlation analysis of data evaluation metrics. (left) We summarize the Pearson correlation coefficients across all six evaluated tasks. (right) We detail the relationship between metric scores (X-axis) and downstream performance (Y-axis) on Language and Code tasks. The Data Verifier maintains a strong, consistent positive correlation across disparate domains. Please refer to Fig. 8 in Appx. D for complete results.
Training. For cold-start SFT, we train Qwen3-32B Yang et al. (2025) on 5K high-quality synthetic instances for 2 epochs, utilizing a learning rate of 2e-5 and a batch size of 32. In the RL phase, we further optimize the SFT checkpoint using GRPO Shao et al. (2024) for 1 epoch on the same dataset, with a learning rate of 5e-7. During RL, the rollout batch size is set to $128$ with a temperature of $1.0$, and we sample 8 candidate data recipes per task.
Evaluation Set. We evaluate on 6 held-out tasks: 3 in-domain tasks and 3 out-of-domain tasks. Notably, these in-domain evaluation tasks share domains with the training set but remain strictly unseen during training. The in-domain benchmarks include PHYSICS Feng et al. (2025), AIME’25 AIME (2025), and LiveCodeBench v6 Jain et al. (2024); the out-of-domain benchmarks are OpenFinData Information (2023), ClimaQA Manivannan et al. (2025), and CHID Zheng et al. (2019).
Metrics. Executing recipes and performing downstream fine-tuning and evaluation are compute-intensive, rendering large-scale end-to-end evaluation impractical. Accordingly, for each evaluation task, we generate a candidate set of $N=32$ independent data recipes. Based on this candidate set, we report two metrics: (1) $\mathrm{DVS}_{\mathrm{avg}@32}$: the mean Data Verifier Score across all 32 recipes. This metric quantifies the expected quality and stability of the policy, where recipes failing to yield valid training data are assigned a score of 0. (2) $\mathrm{DBS}$: the Downstream Benchmark Score of a model trained on a single recipe, which is randomly sampled from the subset of candidates with valid execution (i.e., $\mathrm{DVS}>0$). This metric reflects the actual performance on the downstream benchmark of a successfully executed recipe. Additionally, to approximate the oracle upper bound for DataChef-32B, we select the most promising recipe from the candidate set and report its downstream score. For all downstream evaluation, we fine-tune Qwen3-1.7B-Base for 3 epochs with a learning rate of 2e-5 and a batch size of 64.
Baselines. We compare DataChef-32B against three categories of models: (1) parameter-matched model: Qwen3-32B; (2) open-source flagships: Kimi-K2-Instruct Team et al. (2025) and Qwen3-Next-80B-A3B-Thinking; (3) closed-source SOTA: Gemini-3-Pro Google (2025). Additionally, we incorporate the following results as reference: First, to benchmark the raw data quality, we manually format each available source and report the average ($\textsc{Source}_{\mathrm{avg}}$) and best ($\textsc{Source}_{\mathrm{best}}$) downstream performance among these single-source datasets. Second, as a high-standard reference, we report Expert: the performance of Qwen3-1.7B optimized via industry-grade post-training with expert-curated recipes. Detailed experiment setups are provided in Appx. B.
4.2 Main Results
Main Comparison. Table 1 presents the performance of DataChef-32B against baselines across in-domain and out-of-domain tasks. DataChef-32B achieves superior performance compared to a strong practical baseline, Qwen3-Next $\oplus$ Kimi-K2, which leverages open-source state-of-the-art specialized models (Qwen3-Next-80B-A3B-Thinking for reasoning and Kimi-K2-Instruct for coding). Specifically, our end-to-end model surpasses this composite system with average improvements of +8.6% and +9.2% in $\mathrm{DVS}_{\mathrm{avg}@32}$, and +10.7% and +7.4% in $\mathrm{DBS}$ on in-domain and out-of-domain tasks, respectively. Notably, DataChef-32B achieves performance comparable to the closed-source top-tier Gemini-3-Pro, demonstrating exceptional robustness and effectiveness in automated data recipe generation.
Surpassing Human Baselines. By selecting the most promising recipe from 32 samples (Oracle Upper Bound), DataChef-32B outputforms $\textsc{Source}_{\mathrm{best}}$ on most tasks, achieving an average score of 130.3 on in-domain benchmarks. This indicates that DataChef goes beyond simple dataset selection and synthesizes novel data processing pipelines, including effective selection, mixing, synthesis, and filtering, which are superior to raw manual formatting. Remarkably, it achieves 66.7 on AIME’25 and 46.3 on ClimaQA, even surpassing the Expert baseline with industry-level post-training on expert-curated data recipes. These results underscore the potential of fully automating data recipe generation for LLM training.
Figure 4: Analysis of RL Effectiveness. (a) RL training dynamics indicate that the policy consistently converges toward high-quality data recipe generation. (b) Evaluation results show that RL yields substantial improvements on out-of-domain tasks.
Table 2: Ablation study on training stages and reward design. We investigate the impact of the cold-start phase and the granularity of the reward signal. $\mathcal{R}_{\text{dense}}$ denotes our proposed fine-grained Data Verifier score, while $\mathcal{R}_{\text{sparse}}$ represents a constant success reward for valid execution.
Model Cold Start RL Reward Performance ($\mathrm{DVS}_{\mathrm{avg}@32}$)
In-Domain Out-of-Domain
$\mathcal{M}_{Baseline}$ $\times$ $\times$ - 4.1 5.5
$\mathcal{M}_{RL}$ $\times$ $\checkmark$ $\mathcal{R}_{\text{dense}}$ 32.9 23.9
$\mathcal{M}_{Sparse}$ $\checkmark$ $\checkmark$ $\mathcal{R}_{\text{sparse}}$ 62.7 44.1
DataChef-8B $\checkmark$ $\checkmark$ $\mathcal{R}_{\text{dense}}$ 63.2 46.8
Table 3: Analysis of collaborating with strong coding models. We compare the end-to-end paradigm against decoupled approaches where the model acts solely as a planner, relying on an external coder (Kimi-K2-Instruct) for implementation.
Model External Coder Performance ($\mathrm{DVS}_{\mathrm{avg}@32}$)
In-Domain Out-of-Domain
Inference-Time
Qwen3-32B $\times$ 22.3 18.9
Qwen3-32B 40.3 33.1
Training Paradigm
Planner-32B 56.7 37.3
DataChef-32B $\times$ 64.0 44.1
4.3 Data Verifier
To validate the proposed Data Verifier, we analyze the Pearson correlation between the verifier scores and downstream benchmark performance. We benchmark against several widely used data evaluation metrics, including IFD Li et al. (2024a), RewardModelScore Liu et al. (2025a), DEITA Liu et al. (2024), and VendiScore Friedman and Dieng (2023). To ensure diversity in data quality and model performance, we construct 8–12 datasets per task under a fixed data budget using two strategies: (1) Direct sampling from available task-specific data sources. (2) Subset selection from the pool formed in (1) based on response length. As shown in Fig. 3, with supplementary plots in Appx. D, our Data Verifier exhibits a strong positive correlation with downstream model performance across all six tasks, achieving an average Pearson correlation of $0.59$. Crucially, the Data Verifier maintains consistent positive correlation across diverse task settings, indicating superior robustness compared to baselines.
Figure 5: Analysis of operation frequency in generated recipes. We compare the average number of function calls per recipe across different models.
4.4 Ablation and Analysis
Effectiveness of RL. Fig. 4 illustrates that reward values consistently trend upward while the standard deviation decreases during training, confirming the convergence and effectiveness of RL process. Held-out evaluation reveal that RL primarily enhances generalization, yielding significant gains on out-of-domain tasks while preserving in-domain performance. Quantitatively, RL delivers an average $\mathrm{DVS}_{\mathrm{avg}@32}$ improvement of 3.6% for the 8B model and 3.7% for the 32B model.
Effectiveness of Cold Start. Table. 2 shows that omitting the cold start leads to significant performance degradation across all domains. To understand this behavior, we analyze the distribution of function calls in Fig. 5. The results demonstrate that the direct RL model tends to generate simplistic data pipelines, reducing the usage of complex data processing operations. We hypothesize that without the SFT warm-up, the model succumbs to reward hacking. It avoids execution penalties by generating safe, trivial scripts rather than optimizing for data quality. In contrast, DataChef-8B leverages the SFT foundation to explore and deploy sophisticated operations, such as filtering and data augmentation.
Ablation on Reward Signal. To assess the effectiveness of the fine-grained data quality feedback, we conduct an ablation where the continuous verifier score s(d^)s($\hat${d}) in Eq. 2 is replaced by a constant success reward (i.e., assigning a fixed value of $1.0$ to any valid data recipe). Table 2 demonstrates that this quality-agnostic signal leads to noticeable performance drops. This result confirms that the model relies on the guidance from the Data Verifier to distinguish high-utility recipes from merely executable ones.
Collaborating with Strong Coder. Given the proliferation of specialized coding models, a natural idea is to decouple this task: use the primary model as a planner (natural language orchestration) and an external coder for implementation. Table 3 shows that this paradigm enhances inference-time performance, with Qwen3-32B paired with Kimi-K2-Instruct yielding 18.0% and 14.5% $\mathrm{DVS}_{\mathrm{avg}@32}$ gains on in-domain and out-of-domain tasks, respectively. However, training the model solely as a planner leads to suboptimal results compared to the end-to-end approach. This suggests that integrated training of planning and coding capabilities is essential for optimal data recipe generation.
Figure 6: Visualization of data distribution in generated recipes. We project the source datasets and the data recipes generated by different models into a 2D embedding space.
Case Study. We quantitatively analyze the data recipes generated by Qwen-32B and DataChef-32B for the out-of-domain financial task in Fig 6. We categorize source datasets that yield high downstream performance as High-perf proxy, and those performing poorly as Low-Perf Sources. DataChef-32B demonstrates an emergent ability to identify and prioritize high-utility datasets. Additionally, we provide detailed data processing pipelines with code examples in Appx. C, which reveals that DataChef-32B can: (1) automatically leverage LLMs to augment data into task-specific formats or to synthesize data to enhance target ability; and (2) extract the most relevant data subsets using self-generated keywords.
5 Conclusion
In this paper, we propose a novel paradigm for automated data recipe generation to streamline LLM adaptation. To facilitate this, we establish a holistic dataset for both training and evaluation. Building on this foundation, we present DataChef-32B, incorporating a data verifier that serves as a cost-effective reward function for online RL. DataChef-32B demonstrates strong generalization capabilities, matching human-level expertise on specific benchmarks. Our work bridges the gap between data curation and model evolution, fostering the development of self-evolving AI.
Limitations. Our reliance on an LLM-as-a-Judge for proxy rewards prioritizes generalizability but may sacrifice precision in niche tasks. Developing specialized evaluators to offer higher-resolution reward signals remains a valuable direction for future research.
Appendix A Implementation Details of DataChef
A.1 Details of Task Pool
Table 4: List of benchmarks used in the task pool.
Domain Benchmark Usage
Code HumanEval Chen et al. (2021) Train
LiveCodeBench v6 Jain et al. (2024) Test
Comprehension OpenbookQA Mihaylov et al. (2018) Train
RACE Lai et al. (2017) Train
Instruction Following IFEval Zhou et al. (2023) Train
Agriculture AgXQA Kpodo et al. (2024) Train
SeedBench Ying et al. (2025) Train
Astronomy Astrobench Ting et al. (2024) Train
Biology Genome-Bench Yin et al. (2025) Train
MoleculeQA Lu et al. (2024) Train
ProteinLMBench Shen et al. (2024) Train
Chemistry ChemBench Mirza et al. (2024) Train
Earth Science EarthSE Xu et al. (2025) Train
General Knowledge C-Eval Huang et al. (2023) Train
MMLU Hendrycks et al. (2021) Train
Medical MedXpertQA Zuo et al. (2025) Train
MedQA Yano et al. (2025) Train
Finance OpenFinData Information (2023) Test
Atmosphere ClimaQA Manivannan et al. (2025) Test
Physics PHYBench Qiu et al. (2025) Train
PHYSICS Feng et al. (2025) Test
Long Context L-Eval An et al. (2023) Train
LongBench Bai et al. (2023) Train
Math GSM8K Cobbe et al. (2021) Train
AIME’25 AIME (2025) Test
Reasoning BBH Suzgun et al. (2022) Train
HotpotQA Yang et al. (2018) Train
Safety HaluEval Li et al. (2023) Train
SafetyBench Zhang et al. (2023) Train
Writing WritingBench Wu et al. (2025) Train
Language CHID Zheng et al. (2019) Test
A.2 Prompt Templates and Model Selection
Data Verifier. We employ gpt-oss-120b OpenAI et al. (2025) as the backbone for the Data Verifier. The detailed rubric-based prompt used for evaluation is presented below.
Cold-start Models. To construct high-quality cold-start supervision, we leverage two specialized models: Qwen3-Next-80B-A3B-Thinking Yang et al. (2025) for planning and reasoning, and Kimi-K2-Instruct Team et al. (2025) for code implementation. The specific prompts used for these roles are provided below.
Appendix B Details of Experiments Setup
B.1 Data Evaluation Metrics Settings
We use the OpenDataArena-Tool [^1] https://github.com/OpenDataArena/OpenDataArena-Tool for data assessment, adhering to its default configurations. The specific settings for the data evaluation metrics used in our experiments are as follows:
  1. $\bullet$
    IFD. We employ Qwen2.5-3B-Instruct as the backend model to calculate the Instruction-Following Difficulty (IFD) score. Following Li et al. (2024a), instances with an IFD score >1>1 are treated as outliers. To ensure robust correlation analysis, we assign a score of 0 to these anomalies.
  2. $\bullet$
    DEITA. Following Liu et al. (2024), we define the final data score as the product of the Complexity Score and the Quality Score. These scores are computed using the checkpoints provided in the official DEITA repository.
  3. $\bullet$
    RewardModelScore. We utilize Skywork-Reward-V2-Llama-3.1-8B-40M Liu et al. (2025a) to compute the reward score, serving as a proxy for response quality.
  4. $\bullet$
    VendiScore. We employ Qwen3-Embedding-0.6B to compute sample embeddings and utilize Euclidean distance as the similarity metric to calculate VendiScore, measuring the diversity of the dataset.
B.2 Evaluation Setup
All downstream task evaluations are conducted using the OpenCompass framework [^2] https://github.com/open-compass/opencompass. The detailed settings for each benchmark are as follows:
  1. $\bullet$
    PHYSICS. We employ xVerify-9B-C Chen et al. (2025b) as the evaluator and report the average accuracy across all sub-tasks.
  2. $\bullet$
    AIME’25. We evaluate on the 2025 subset (covering both Part I and Part II). For each question, we generate 8 responses and report the average accuracy. xVerify-9B-C is used as the evaluator.
  3. $\bullet$
    LiveCodeBench v6. We utilize the official prompt guidelines and report the pass@1 metric. The LCBCGenerationEvaluator is used for assessment.
  4. $\bullet$
    ClimaQA. We employ xVerify-9B-C as the evaluator and report the average accuracy across all sub-tasks.
  5. $\bullet$
    OpenFinData. We use the OpenFinDataKWEvaluator and report the average accuracy across all sub-tasks.
  6. $\bullet$
    CHID. We report the average accuracy on both the development and test sets.
Prompt for Data Recipe Generation (Natural Language) # Task Description
{{ task_description }}
# Benchmark
## {{ benchmark.name }}
{{ benchmark.description }}
# Available Hugging Face Training Datasets
{% for item in datasets -%}
## {{ item.dataset_id }}
{{ item.examples }}
{% endfor -%}
Based on the Task Description, the target Benchmark, and the Available Hugging Face Training Datasets, design a feasible Data Processing Plan. This plan must include: (1) **Data Selection**: Identify the most suitable raw datasets from the available list. (2) **Data Processing Workflow**: Define the pipeline to transform selected raw data into high-quality SFT training data. The plan will serve as a blueprint for code generation and data production. Ensure it is comprehensive, actionable, and free from ambiguous or vague statements. A high-quality SFT dataset must exhibit the following attributes: High Quality: Accurate samples free from noise. Diversity: Coverage of varied instructions and objectives. Relevance: Strong alignment with the target vertical domain. Important Constraints & Guidelines: Source Selection: strictly select datasets provided in the context. Prioritize those with complete metadata and clear field definitions. Strictly prohibit data contamination: DO NOT use benchmark data for training. Do not hallucinate datasets, splits, or configurations not provided in the context. Grounding: Base the processing workflow solely on the actual fields and content of the selected datasets. Do not make assumptions about data fields that do not exist. Context: Ensure inputs and outputs form coherent Q&A turns. If the output relies on specific context (e.g., a document or snippet) within the data, design the workflow to explicitly embed this context into the input. Format Alignment: If the benchmark requires specific formats (e.g., multiple-choice, JSON output), ensure the constructed training data aligns with these requirements. Final Output Format: The pipeline must produce data in the following standard dialogue format:
{"dialogs": [{"role": "user", "content": "..."}, {"role": "assistant", "content": "..."}]}
LLM Utilization: Flexibly leverage LLM inference capabilities for tasks such as: extracting instructions from heterogeneous documents, data augmentation/synthesis, and quality verification. Clearly specify the prompt design strategies for these steps.
Output Format: Generate the plan strictly in the following format (do not include preamble or additional text): ## Training Data
{"dataset_id": "target_id", "split": "target_split", "name": "config_name", "sample_num": "int", "reason": "justification"},
## Data Processing Workflow
[Detailed Data Processing Workflow description]
Prompt for Data Recipe Generation (Executable Code) # Available Hugging Face Training Datasets
{% for item in datasets -%}
## {{ item.dataset_id }}
{{ item.examples }}
{% endfor -%}
# Data Processing Plan
{{ plan }}
# Tool Information
{{ tool_info }}
Based on the Available Hugging Face Training Datasets, the Data Processing Plan, and the Tool Information, generate the executable **Data Processing Script**. To validate the correctness of the processing, also generate a corresponding **Verification Script**. Important Notes & Constraints: Implementation Logic: The script must clearly implement the processing rationale and utilize tools correctly. Verification Scope: The verification script should validate that the generated data matches expectations and confirm the effectiveness of core processing steps. Target Format: The final data must follow the ShareGPT format:
{’dialogs’: [’role’: ’user’, ’content’: ’...’, ’role’: ’assistant’, ’content’: ’...’]}. Use the format_to_sharegpt utility (import from aidp).
Output Directory: Save the processed data to the data/processed/ directory.
Output Format: Generate exactly two Python code blocks in the following format, without any additional text: ‘‘‘python
# data-processing code block
‘‘‘python
# test code block
[…]
Appendix C Case Study
To demonstrate the capability of our model, we present a complete data processing pipeline generated by DataChef-32B for the ClimaQA task. As shown in Fig. 7, the generated code successfully produces valid training data by: (1) automatically leveraging LLMs to augment data into task-specific formats and synthesize samples to enhance target capabilities; and (2) extracting the most relevant data subsets using self-generated keywords.
1
2# data-processing code block
3...
4
5# 1. Load stackexchange_earthscience
6stack_ds = load_remote_dataset(
7 "mlfoundations-dev/stackexchange_earthscience",
8 name="default",
9 split="train",
10 shuffle=False
11)
12
13# 1.1 Use LLM to convert Open QA into 4-option Multiple Choice Questions
14SYSTEM_PROMPT_STACK = """
15You are a science tutor.
16Convert the given open QA pair into a 4-option multiple-choice question.
17Return JSON only:
18{"question": "<original question>", "A": "<option A>", "B": "<option B>", "C": "<option C>", "D": "<option D>", "answer": "<correct letter>"}
19The correct answer must be one of the four options. Distractors should be scientifically plausible but wrong.
20"""
21
22def stack_parser(response: str, raw: dict) -> list:
23 parsed = extract_json(response)
24 if not parsed or "answer" not in parsed:
25 return []
[…]
28 assistant = parsed["answer"].lower()
29 return [{"user": user, "assistant": assistant}]
30
31stack_mc = generate_dataset_with_llm(
32 stack_ds,
33 system_prompt=SYSTEM_PROMPT_STACK,
34 response_parser=stack_parser
35)
Generated Data Recipe Code (ClimaQA Case Study Part I)
1
2# 2. Load ScienceQA_text_only
3sciqa_ds = load_remote_dataset(
4 "tasksource/ScienceQA_text_only",
5 name="default",
6 split="train",
7 shuffle=False
8)
9
10# 2.1 Filter for climate-related topics
11def climate_filter(sample):
12 txt = (sample.get("category", "") + " " + sample.get("topic", "")).lower()
13 return any(k in txt for k in ("climate", "weather", "atmosphere", "ocean", "environment"))
14
15sciqa_climate = select_by_filter(sciqa_ds, climate_filter)
16
17# 2.2 Standardize format
18def sciqa_map(sample):
19 choices = sample["choices"]
[…]
22 assistant = chr(65 + sample["answer"])
23 return {"user": user, "assistant": assistant}
24
25sciqa_mc = sciqa_climate.map(sciqa_map, remove_columns=sciqa_climate.column_names)
26
27# 3. Merge & Deduplicate
28merged = concatenate_datasets([stack_mc, sciqa_mc])
29deduped = deduplicate_by_text_hash(
30 merged,
31 text_map=lambda x: x["user"],
32 lowercase=True,
33 ignore_non_character=True
34)
35
36# 4. Convert to ShareGPT format
37sharegpt_ds = format_to_sharegpt(
38 deduped,
39 user_map=lambda x: x["user"],
40 assistant_map=lambda x: x["assistant"]
41)
42
43# 5. Save output
44dump_dataset(sharegpt_ds, "data/processed/train_climaqa_style.jsonl")
Generated Data Recipe Code (ClimaQA Case Study Part II)
Figure 7: Case study of data recipe generation.
Appendix D Additional Results on Correlation Analysis
We provide the comprehensive correlation analysis results across all six evaluation tasks in Fig. 8. These results validate the robustness of our Data Verifier compared to baseline metrics across diverse domains.
Figure 8: Complete results for correlation analysis.